OpenStudy (anonymous):

The sum of two numbers is \(\color{green}{6}\) times their \(\color{blue}{Geometric Mean}\). Show that the Numbers are in the Ratio - \(\color{red}{(3+2\sqrt{2}) : (3 - 2 \sqrt{2})}\)

OpenStudy (accessdenied):

2\{sqrt dangerous {

OpenStudy (anonymous):

Oh man where is the problem..

OpenStudy (accessdenied):

Oh, looks good now! :)

OpenStudy (anonymous):

Oh sorry forgot to refresh the page,..

OpenStudy (anonymous):

Let us come to the question now...

OpenStudy (anonymous):

\[a+b = 6 \sqrt{ab}\]

OpenStudy (anonymous):

\[\frac{a}{b} = \frac{6 \sqrt{a}}{\sqrt{b}} - 1\]

hero (hero):

\[\frac{a}{b} = \frac{(3+2\sqrt{2})}{ (3 - 2 \sqrt{2})}\]

OpenStudy (anonymous):

This, I think we have to prove..

hero (hero):

I know

OpenStudy (anonymous):

Yeah, I know too...

hero (hero):

I'm pretty sure you can figure it out. Its just algebraic manipulation.

OpenStudy (anonymous):

Let me try then..

OpenStudy (anonymous):

I am not getting where to start from.. Should I divide both the sides by sqrt{ab} ??

hero (hero):

I figured it out without doing any paperwork

OpenStudy (anonymous):

That is why you are Hero I think..

OpenStudy (anonymous):

\[\frac{a}{b} = 17 + 12 \sqrt{2}\]

hero (hero):

?

OpenStudy (anonymous):

this is some more simplification of what we have to prove.. Sorry not we, but me..

hero (hero):

I think there is an easier way bro

OpenStudy (anonymous):

Wait, I am thinking then... Yeah there must be..

hero (hero):

Just let me know when you give up

OpenStudy (anonymous):

Sure why not..

hero (hero):

Did you figure it out yet?

OpenStudy (anonymous):

No, I am still trying it on my notebook..

OpenStudy (anonymous):

Sorry, internet got disconnected..

OpenStudy (anonymous):

Give me some hint, I am not getting where to start from..

OpenStudy (anonymous):

Give me some hint if you can, I am not getting where to start from..

OpenStudy (shubhamsrg):

well this ain;t much of the problem.. you already know a/b = 6sqrt(a/b) -1 just let sqrt(a/b) = x.. hope this helps..

OpenStudy (anonymous):

I am hoping too.. Wait then, I will go forward now..

OpenStudy (anonymous):

\[x = 3 \pm 2 \sqrt{2}\]

OpenStudy (anonymous):

\[\frac{\sqrt{a}}{\sqrt{b}} = 3 \pm 2 \sqrt{2}\]

OpenStudy (anonymous):

@hero Am I right??

OpenStudy (anonymous):

\[\frac{\sqrt{a}}{\sqrt{b}} = 3 + 2 \sqrt{2}\] \[\frac{\sqrt{a}}{\sqrt{b}} = 3 - 2 \sqrt{2}\]

OpenStudy (anonymous):

If you let:\[a+b=k\]then it follows that:\[6\sqrt{ab}=k\Longrightarrow ab=\frac{k^2}{36}\]Therefore you want to solve this system:\[a+b=k\]\[ab=\frac{k^2}{36}\]which is the same as solving the quadratic:\[x^2-kx+\frac{k^2}{36}=0\] you can use the quadratic formula formula from here and divide the answers to get the desired ratio.

OpenStudy (anonymous):

Wait I am finding here x as you are saying...

OpenStudy (anonymous):

\[x = \frac{k \pm \sqrt{k^2 - \frac{k^2}{9}}}{2}\]

OpenStudy (anonymous):

thats right :) so let a be the solution with the +, and b be the solution with the -. Clean it up a little be too. You can factor the k^2 inside the radical, etc etc.

OpenStudy (anonymous):

Yeah I am, Net is very slow..

OpenStudy (anonymous):

\[x = \frac{3k \pm 2k \sqrt{2}}{2}\]

OpenStudy (anonymous):

Wait there is 6 in the denominator...

OpenStudy (anonymous):

nice, so if:\[a=\frac{3k+2k\sqrt{2}}{2}\]and\[b=\frac{3k-2k\sqrt{2}}{2}\] then:\[\frac{a}{b}=\]

OpenStudy (anonymous):

I am not getting this that how you have taken a and b from here??

OpenStudy (anonymous):

Ah, im sorry, i should have explained it better. So why is solving the system:\[a+b=k\]\[ab=\frac{k^2}{36}\]the same as solving the equation:\[x^2-kx+\frac{k^2}{36}=0\]? Its because when you are solving the quadratic equation:\[x^2-bx+c=0\]by factoring, you ask yourself, "What are two numbers that add up to b, and multiply to c?" But this statement is exactly the system:\[m+n=b\]\[mn=c\] basically, solving the system of equations, or just the quadratic yield the same solutions. That is why im saying the solutions to the quadratic are now a and b.

OpenStudy (anonymous):

Oh so a and b are the roots of the equation...

OpenStudy (anonymous):

exactly :)

OpenStudy (anonymous):

I did not look that way.. Sorry I got it.. Thanks you for explaining me better in this regard...

OpenStudy (anonymous):

*Thank.. And also thanks to @Hero Sorry I could not follow you...